Suppose that a guitar string has a length of 1 m, a mass of 0.15 kg, and a tension of 140 N. Find:

a) The mass per unit of length of this string

b) The speed of a wave on this string


1
Expert's answer
2021-09-12T19:01:43-0400

a) We can find the mass per unit of length of this string as follows:


"\mu=\dfrac{0.15\ kg}{1\ m}=0.15\ \dfrac{kg}{m}."

b) We can find the speed of a wave on this string as follows:


"v=\sqrt{\dfrac{T}{\mu}}=\sqrt{\dfrac{140\ N}{0.15\ \dfrac{kg}{m}}}=30.5\ \dfrac{m}{s}."